OpenStudy (lgbasallote):

Prove that if n is a positive integer then n is even if and only if 7n + 4 is even.

OpenStudy (lgbasallote):

What is the solution for iff?

OpenStudy (mayankdevnani):

If n is even

OpenStudy (mayankdevnani):

then there is an integer k with n = 2k.

OpenStudy (mayankdevnani):

ok

OpenStudy (lgbasallote):

i was asking about iff

OpenStudy (lgbasallote):

not the solution...yet

OpenStudy (mayankdevnani):

no it is right

OpenStudy (lgbasallote):

huh?

OpenStudy (anonymous):

first of all lets take, if n is even then then the answer of 7n+4 will be even only

OpenStudy (lgbasallote):

first of all...what is iff?

OpenStudy (mayankdevnani):

7n+4=7(2k)+4 14k+4=2(7k+2)

OpenStudy (mayankdevnani):

ok

OpenStudy (anonymous):

if you multiply 7 with any even number you will get the ans as even number only.that is the very valid reason for this.

OpenStudy (lgbasallote):

im asking about iff not the solution @mayankdevnani

OpenStudy (lgbasallote):

im not asking about the solution to this problem.. im just asking what iff is

OpenStudy (anonymous):

and adding 4 will does not make any difference to number type whether it is "odd" or "even".

OpenStudy (lgbasallote):

@rambo2210 are you answering my question?

OpenStudy (anonymous):

"iff" ?? where is it?

OpenStudy (lgbasallote):

if and only if

OpenStudy (lgbasallote):

im just asking what it means

OpenStudy (anonymous):

@lgbasallote ok let me explain you "iff"..

OpenStudy (lgbasallote):

sure

OpenStudy (anonymous):

iff means "if and only if".. it means if "7n+4" will answer the even number then "n" will be even for sure.else it will be odd.

OpenStudy (anonymous):

Now you got it?

OpenStudy (lgbasallote):

wait....let me process it

OpenStudy (lgbasallote):

so how is it different from if-then?

OpenStudy (anonymous):

If-then is conditional

OpenStudy (lgbasallote):

what about in terms of solution...how is it different from if-then?

OpenStudy (anonymous):

post you if-then question here for example.

OpenStudy (lgbasallote):

i think im seeing the difference in wording now (there are two conditions in this case) but anyway... p = n is even q = 7n + 4 is even assume n is even n = 2k 7n + 4 7(2k) + 4 14k + 4 2(7k + 2) let 7k + 2 = x 2x \(\therefore\) 7n + 4 is even when n is even

OpenStudy (lgbasallote):

so how to solve it using biconditional?

OpenStudy (anonymous):

what is bi-conditional?

OpenStudy (lgbasallote):

that's what i want to know too

OpenStudy (anonymous):

the logical biconditional (sometimes known as the material biconditional) is the logical connective of two statements asserting "p if and only if q", where q is a hypothesis (or antecedent) and p is a conclusion (or consequent).[1] This is often abbreviated p iff q.

OpenStudy (lgbasallote):

well i know that part....what about the solution?

ganeshie8 (ganeshie8):

a iff b => if a, then b AND if b, then a

ganeshie8 (ganeshie8):

if we prove both directions, we're done

OpenStudy (lgbasallote):

ahh so you prove p -> q and q -> p?

ganeshie8 (ganeshie8):

i think so

OpenStudy (lgbasallote):

wonderful

OpenStudy (anonymous):

yeap ganeshie8 is rgt.